OpenStudy (anonymous):

If you wanted to eliminate the y variable by adding the two equations in the following system, what could be your first step? x − y = 6 4x − 3y = 12

OpenStudy (zepp):

You have to make a 3y appear in your first equation, or, make the 3 disappear in the second equation.

OpenStudy (zepp):

\[\frac{4x − 3y = 12}{3}\rightarrow\frac{4}{3}x-y=4\]

OpenStudy (zepp):

Or, (x − y = 6)3 \(\rightarrow\) 3x-3y=18

OpenStudy (anonymous):

Add the two equations together Multiply the first equation by −4 Multiply the first equation by 3 Multiply the first equation by −3

OpenStudy (zepp):

Oh, adding up. Then (x − y = 6)-3=-3x+3y=-18

OpenStudy (zepp):